The perimeter of a rectangle is 46 feet. The width is 3ft less than twice the length. Find the length and width of the rectangle?
Answer by oscargut(2103)

Solution:
Let L= length and W =width
W = 2L-3
Perimeter = 2(L+W)=2(L+2L-3) = 2(3L-3)=6L-6 = 46
6L = 46 + 6
6L = 52
L = 52/6
L = 26/3
W = 2(26/3)-3 = 52/3-9/3 = 43/3
Answer: Width = 43/3 and Length = 26/3
You can send me more problems to my e-mail: mthman@gmail.com